Find the nth term of the geometric sequence.
0.3, 0.03, 0.003, ...

Answer by stanbon(75887)

Find the nth term of the geometric sequence.
0.3, 0.03, 0.003, ...
----
a1 = 3/10
a2 = 3/10^2
a3 = 3/10^3
------
an = 3/10^n
=================
Cheers,
Stan H.